# Zens, Díaz, Durante and Patacchini

### v1

@@ Abstract
The presentation of the results of the analysis under the proposed model is divided in four parts. In Section 5.2.1, we first compare the forecasting accuracy of model (1)–(2) with those achieved by a full-rank bilinear autoregressive formulation and a black-box predictive strategy relying on random forests. Besides clarifying that the proposed parsimonious representation is competitive with strategies specifically designed for forecasting, this assessment provides reassurance on the fact that inference on the interpretable dependence structures encoded in $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, ${\bf A}$ and ${\bf B}$ relies on a realistic characterization of the observed criminal activity tensor. The results of inference on the posterior point estimates for these quantities are discussed in Sections 5.2.2 and 5.2.3, respectively. Finally, in Section 5.2.4 we interpret the findings arising from such inferences in the light of the general qualitative and quantitative evidence on criminal activities presented within Section 2.
All the results presented are based on $3{,}000$ posterior samples for the quantities of interest produced by Algorithm 1, following an initial burn-in period of $1{,}000$ iterations. Convergence and mixing of such an algorithm posed no major challenges. To select the ranks $R_{A}$ and $R_{B}$, we conducted an exhaustive enumeration of all possible values for the pair $(R_{A},R_{B})$, evaluating these combinations in terms of out-of-sample predictive accuracy, similarly to the assessments conducted in Section 5.2.1. The results of this analysis suggest $R_{A}=16$ and $R_{B}=8$. This is the lowest rank specification for the proposed model (1)–(2) that yields an out-of-sample predictive accuracy within $1\%$ of the one achieved by a fully-rank bilinear vector autoregressive formulation.
@@
Before discussing the results of inference under model (1)–(2), let us first assess its forecasting performance and compare it against alternative predictive strategies. In presenting the results of such an analysis, it is important to emphasize that while a significant portion of the empirical literature on crime data analysis is primarily designed for purely-predictive purposes, the model developed in Sections 34 aims instead at providing interpretable inference on the core dependence structures within the observed criminal activity tensor. As such, the overarching goal of our predictive assessment and comparisons is not to demonstrate that the proposed model systematically outperforms all potential competitors, including black-box ones, in out-of-sample forecasting. Rather, we aim to provide empirical evidence on the fact that carefully-designed model-based representations, such as the one we developed, can substantially enhance interpretability and inference potentials, while preserving a predictive accuracy that is competitive with those achieved by more sophisticated strategies, including highly-parameterized formulations and black-box solutions. Such an assessment also serves as a sanity check to quantify whether the proposed model yields a realistic representation of the generative mechanisms underlying the data analyzed, thereby reassuring on the reliability of the inference results in Sections 5.2.2 and 5.2.3.
mse mae corr
[…]
Full-rank bilinear autoregression 0.380 0.279 0.735
Random forests 0.367 0.272 0.755
[…]
@@
Table 1: Forecasting performance of the proposed model (low-rank bilinear autoregression) and relevant competitors (full-rank bilinear autoregression, random forests) under different metrics: mse (mean squared error), mae (mean absolute error) and corr (correlation between predictions and actual observed values). Results are averaged across 24 replicated one-step-ahead forecasting exercises. The proposed low-rank bilinear autoregressive model employs $R_{A}=16$, $R_{B}=8$.
@@
Consistent with the above goals, the predictive assessment is conducted by dividing the data into a training set and a test set. The training sets cover data ${\bf Y}_{1},\ldots,{\bf Y}_{t_{\mbox{\tiny train}}}$ for 204 consecutive months. Leveraging these training data we fit the competing methods under analysis, including the proposed model, and then produce one-step-ahead forecasts for the reported criminal activities at the time $t_{\mbox{\tiny train}}+1$ for all the 616 combinations of districts and crime categories. The resulting forecasts are finally compared to the actually observed values in ${\bf Y}_{t_{\mbox{\tiny train}}+1}$ using three different accuracy measures, namely the mean squared error (mse), the mean absolute deviation (mae) and the correlation (corr) between the forecasted values and actual observed ones in ${\bf Y}_{t_{\mbox{\tiny train}}+1}$. To provide a comprehensive assessment, these measures are further averaged over 24 replications of such a predictive analysis, each obtained by shifting the training window of 204 observations forward by one month, and forecasting the criminal activities in the subsequent month. As a result, predictive performance is quantified on a total of $616\cdot 24=14{,}784$ point forecasts.
Table 1 summarizes the outcome of the aforementioned predictive study for the proposed model (low-rank bilinear autoregression with $R_{A}=16$, $R_{B}=8$) and two relevant competing methods. The first one (full-rank bilinear autoregression) still relies on the bilinear representation as in (1), but does not include the low-rank structure in (2). The second (random forests (Breiman, 2001)) provides instead a relevant example of a routinely-implemented black-box strategy specifically designed for predictive purposes. As such, one would expect that this latter strategy would substantially improve the forecasting accuracy of the low-rank and full-rank bilinear autoregressive models. As illustrated in Table 1, this is not the case in such a context. In fact, both model-based solutions achieve a predictive accuracy competitive with random forests, with the low-rank formulation (1)–(2) providing essentially the same performance as its full-rank counterpart. These results clarify that interpretable model-based inference can be accomplished at negligible costs in predictive accuracy, and provides reassurance that the results presented within Sections 5.2.25.2.3 are reliable and not based on an unrealistic low-rank characterization of the data-generating mechanism.
@@
Besides achieving satisfactory predictive accuracy, a distinctive feature of the model developed in Section 34 lies in its ability to disentangle complex dependencies in the criminal activity tensor, while allowing interpretable inference on such quantities. To showcase such potentials, let us first focus on the intra-temporal dependence structures among crime categories and between districts encoded in the posteriors of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$. Figure 3 provides a graphical representation of these estimated dependences with a focus on the posterior mean of the partial correlation networks defined as $-\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}\bm{\Sigma}_{A}^{-1}\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}$ and $-\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}\bm{\Sigma}_{B}^{-1}\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}$ (e.g., Lauritzen, 1996). Such posterior means can be obtained by evaluating these formulas at the posterior samples of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ produced by Algorithm 1 and then averaging across samples. Crucially, these measures provide an informative reconstruction of the graphical dependencies between crime categories and across the different districts, which quantify the intra-temporal pairwise correlations among any two generic nodes in the network conditionally on the effect of all the others, and net of the inter-temporal dependencies captured by the conditional mean in (1)–(2).
Figure 3: Estimated intra-temporal partial correlation networks between crime categories (left) and among districts (right). The colors in each network indicate clusters of nodes with high within-group absolute partial correlations and relatively-low across-group dependencies (the clusters are estimated via community detection algorithms (Blondel et al., 2008)). Node positions are obtained via force-direct placement methods (Fruchterman and Reingold, 1991) positioning closer in the space nodes with higher absolute partial correlations.
Figure 4: Left: Districts of Chicago colored according to the communities estimated from the partial correlation network in the left panel of Figure 3. Right: Plot of the log-transformed distances between pairs of district centroids versus the posterior mean of the partial correlation among such district pairs (loess estimate and uncertainty bands for the relation between these two quantities are also provided).
@@
Figure 3 represents graphically the aforementioned networks with nodes corresponding to crime categories and districts, respectively, and weighted edges denoting the posterior mean of the partial correlation among any pair of nodes. To facilitate interpretation and uncover relevant structures, the positions of the nodes within Figure 3 are obtained via force-direct placement methods (Fruchterman and Reingold, 1991), so that nodes with higher absolute partial correlations are displayed closer in the space. This graphical setting is also combined with the output of a community detection algorithm (Blondel et al., 2008) that unveils clusters of nodes having high within-group absolute partial correlations and relatively-low across-group dependencies. As shown in the network among the different crime categories, such a perspective allows us to uncover yet-unexplored structured modular intra-temporal dependencies between types of crimes. More specifically, the first group (green nodes) comprises crimes such as simple and aggravated battery and assault, which are physical crimes directly targeting the body. The second (purple nodes) includes, instead, crimes related to narcotics, gambling, weapons violations, and liquor license-related offenses. This clustering suggests a potential association with organized crime or coordinated gang activities in the city. The third group (blue nodes) consists mainly of property crimes such as burglary, robbery, theft, vandalism, and trespassing. These types of crime are often interconnected for various reasons. For example, the evidence suggests that the presence of vandalism, like graffiti, can significantly increase the propensity of people to engage in property crimes, such as stealing (Keizer et al., 2008). At the same time, it is worth noting that the connection could be directly related to legal definitions. To give an instance, trespassing evolves into burglary when an individual trespasses with the intent of committing a crime during that trespass (i.e., the difference pertains to whether the crime in question was attempted or actually committed). Finally, the last group (red nodes) captures crimes which exhibit sparse correlations, even within the community itself. This cluster encompasses crimes such as arson, homicide or kidnapping, which are offenses that tend to occur independently of other crimes. A partial exception in such a group is provided by prostitution, which appears in Figure 3 relatively close to the purple nodes, thereby suggesting that this category might have also direct connections with organized crime or coordinated gang activities. See Section 5.2.4 for further discussion on these findings in the light of the qualitative and quantitative criminology theories presented in Section 2.
@@
Similarly-interesting results can be obtained by examining the network structure of the partial correlations among the districts displayed in the right panel of Figure 3. The community detection algorithm (Blondel et al., 2008) applied to such a network reveals in this case two distinct groups, which fully align with the north-south division of the city of Chicago. This geographical separation becomes even more evident when visualizing the estimated group structure on a map of the city, as demonstrated in the left panel of Figure 4. The connection of the partial correlations to geographical distance is further underlined in the right panel of Figure 4, which highlights a negative relationship between the pairwise distances among the districts and the corresponding partial correlation. As discussed in Section 5.2.4 these findings contribute to the ongoing studies on criminal activities within urban environments and provide also useful resources to quantify potential spillover effects across the districts. Remarkably, the proposed formulation identifies these spatial dependence structures without including any explicit geographical information within the model. Although it would be interesting to extend the proposed framework to explicitly incorporate this additional source information, such an extension is beyond the scope of the current article and might even undermine the flexibility of the model. In fact, the spatial dependencies that we learn might arise from the direct effect of underlying latent factors such as localized activities of criminal organizations or distinctive policing strategies, which are often difficult to include in the model only through geographical information. In fact, avoiding such constraints allows us to learn in the left panel of Figure 4 stronger dependencies among southern districts relative to those inferred for northern ones. This result partially departs from pure geographical proximity and provides important quantitative evidence on the presence of organized crime and structured gangs activities in the southern area of Chicago (e.g., Aspholm, 2019).
@@
Figure 5 displays the heatmaps of the posterior means for the elements in the matrices ${\bf A}$ and ${\bf B}$. Recalling Section 3.1, these quantities provide insights into the average inter-temporal dependence among the $K$ crime categories within districts, and between the $Q$ districts for each crime category, respectively. As such, although these estimates may not directly correspond to causally-interpretable effects, the results in Figure 5 are still able to unveil relevant association patterns. First, the dominant feature of both matrices is the main diagonal, indicating a moderately-high inter-temporal persistence within reports by crime categories and districts, respectively. In the posterior mean of the matrix ${\bf A}$ (see left panel of Figure 5) certain crimes, such as arson and homicide, occur infrequently and in relative isolation, leading to exceptions in this overall pattern. Additionally, the occurrence of specific crime categories, such as simple battery, vandalism, and theft-related crimes, appears to have a more remarkable association with the potential outbreak of specific types of crimes in the subsequent month. Such a finding provides important guidelines in the design of intervention policies guided by potential spillover effects.
Similarly interesting results are provided by the posterior mean of the matrix ${\bf B}$ in the right panel of Figure 5. Besides the direct inter-temporal persistence highlighted in the main diagonal, the estimated ${\bf B}$ clarifies that such a persistence acts also at a group level, especially for clusters of southern districts whose crime activities appear to be associated with those at subsequent times in other districts from the same group. This result is interestingly in line with the spatial intra-temporal dependencies discussed in Section 5.2.2 and further support the idea that these patterns might arise from the presence of organized crime and coordinated gangs activities in the southern area of the city (e.g., Aspholm, 2019).
Figure 5: Heatmaps of the posterior mean for the matrices ${\bf A}$ (left) and ${\bf B}$ (right) characterizing the inter-temporal dependencies under model (1)–(2).
@@
The results in Sections 5.2.25.2.3 provide novel empirical evidence in the context of criminological analyses on the mechanisms and determinants driving concentration of specific crime types in time and space (e.g., Brantingham and Brantingham, 2016, 2017; Cohen and Felson, 1979). Social, economic, and even structural characteristics of urban environments influence criminal behavior (Brantingham and Brantingham, 1995), leading to specific inter and intra-temporal dependencies within both crime categories and spatial districts that also exhibit peculiar modular patterns (see Figure 3 and Figure 5). As such, our results further underscore the importance of maintaining a systemic perspective on crime phenomena, accounting for all the three dimensions of the observed criminal activity tensor, while inferring interpretable dependencies within and across such dimensions. For example, the spatial effects we unveil in Figure 3 and Figure 5 highlight an heterogeneous distribution of crime generators and crime attractors within the urban environment (Brantingham and Brantingham, 1995; Weisburd, 2015), which, in turn, do not facilitate the same crime categories across the whole city. An interesting case is provided by the Chicago Loop located in the 1st district. This area serves as a financial, cultural, and commercial hub of the city, thereby increasing the number of potential targets, especially in relation to property crime. The right panels of Figure 3 and Figure 5 show that, although the 1st district is located in the northern group, it still provides a crime generator area that showcases non-negligible intra- and inter-temporal associations with southern districts. Recalling, e.g., Aspholm (2019), these southern districts tend to act, instead, as crime attractors due to reduced interventions and structural organizational issues that create opportunities for organized criminal activities, and hinder residents from exerting informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942). North Lawndale, situated in the 10th district (southern group in Figure 3 and Figure 4), exemplifies this dynamic. As one of the most impoverished and socially disorganized neighborhoods in Chicago, it experiences high rates of homicides and gun violence. These criminogenic disparities across the different areas of Chicago seem to emerge in our findings, which highlight clear modular separations in the intra- and inter-temporal associations within and across northern and southern Chicago districts, with the latter exhibiting stronger within-group dependencies.
@@
The results within Figure 4 further underscore the importance of the geographic proximity and potential spillover effects. According to our findings, there is a negative relationship between the distance separating pairs of districts and the corresponding partial correlation. One possible explanation involves mechanisms that combine residential location and social interactions (Zenou, 2003), highlighting the role of behavior in disadvantaged environments (Glaeser et al., 1996). For instance, evidence suggests that offenders are inclined to exploit criminal opportunities close to their residential areas (e.g., Bernasco, 2010; Bernasco and Block, 2009; Bernasco et al., 2014; Hipp, 2016; Menting et al., 2016; Townsley et al., 2016). Consequently, the concentration of potential offenders within a neighborhood can trigger a contagion effect, propagating criminal behavior and reinforcing negative social dynamics (Becker and Murphy, 2009).
@@
Moreover, increased criminal concentration intensifies the competition among offenders, reducing the expected gain of criminal activity (Becker, 1968) and encouraging repeat offenses (Zenou, 2003). This is also consistent with research on repeat victimization and temporal clustering, which indicates that the occurrence of an initial crime increases the probability of subsequent offenses in the same location. Offenders tend to exploit previously-acquired knowledge about susceptible targets and the absence of effective interventions (e.g., Johnson and Bowers, 2004; Townsley et al., 2003). This persistence of crime is also evident in our findings (see Figure 5), although there are differences across crime categories. Specifically, while the main diagonal of the left matrix within Figure 5 shows a clear temporal persistence for property crimes such as theft, robbery, and burglary, as well as for organized crime-related offenses like narcotics, prostitution, and gambling, this pattern is less pronounced for the violent offenses such as sexual-related crimes, homicides, and manslaughter, and for the property crimes like arson and embezzlement. These inferred heterogeneities in crime persistence by offense type provide fundamental empirical evidence to the literature on such phenomena (Chainey, 2021; Haberman and Ratcliffe, 2012; Johnson and Bowers, 2004; Johnson et al., 2009).

### v2

@@ Abstract
The presentation of the results of the analysis under the proposed model is divided in four parts. In Section 5.2.1, we first compare the forecasting accuracy of model (1)–(2) with those achieved by a full-rank bilinear autoregressive formulation and a black-box predictive strategy relying on random forests. Besides clarifying that the proposed parsimonious representation is competitive with strategies specifically designed for forecasting, this assessment provides reassurance on the fact that inference on the interpretable dependence structures encoded in $\bm{\Sigma}_{A}$, $\bm{\Sigma}_{B}$, $\bf A$} and $\bf B$} relies on a realistic characterization of the observed criminal activity tensor. The results of inference on the posterior point estimates for these quantities are discussed in Sections 5.2.2 and 5.2.3, respectively. Finally, in Section 5.2.4 we interpret the findings arising from such inferences in the light of the qualitative and quantitative evidence on criminal activities presented in Section 2.
All the results presented leverage Monte Carlo estimates based on 3,0003{,}000 posterior samples for the quantities of interest produced by Algorithm 1, following an initial burn-in period of 1,0001{,}000. The total runtime of this routine is 20 minutes on a standard laptop, much lower than the monthly time scale at which data are analyzed, and hence, the estimates would require updating. Convergence and mixing of such an algorithm posed no major challenges. To select the ranks $R_{A}$ and $R_{B}$, we conducted an exhaustive enumeration of all the possible values for $(R_{A},R_{B})$, evaluating these combinations in terms of out-of-sample predictive accuracy, similarly to the assessments conducted in Section 5.2.1. The results of this analysis suggest $R_{A}=11$ and $R_{B}=5$. This is the lowest rank specification for the proposed model (1)–(2) that yields an out-of-sample predictive accuracy within $0.5\%$ of the one achieved by a full-rank bilinear vector autoregressive formulation.
@@
Before discussing the results of inference under model (1)–(2), let us first assess its forecasting performance and compare it against alternative predictive strategies. In presenting the results of this analysis, it is important to emphasize that while a significant portion of the literature on crime data analysis is primarily designed for purely predictive purposes, the model developed in Sections 3 and 4 aims instead at providing interpretable inference on the dependence structures within the observed criminal activity tensor. As such, the overarching goal of our predictive assessment is not to demonstrate that the proposed model systematically outperforms all potential competitors, including black-box ones, in out-of-sample forecasting. Rather, we aim to provide empirical evidence on the fact that carefully designed model-based representations, such as the proposed one, can substantially enhance interpretability and inference potentials, while preserving a predictive accuracy that is competitive with those achieved by highly parameterized formulations and black-box solutions. This assessment also serves as a sanity check to quantify whether the proposed model yields a realistic representation of the generative mechanism underlying the data analyzed, thereby reassuring on the reliability of the inference results in Sections 5.2.25.2.3.
rmse mae corr
Low-rank bilinear autoregression 0.375 0.270 0.974
Full-rank bilinear autoregression 0.373 0.269 0.975
Random forests 0.365 0.270 0.976
@@
Table 1: Forecasting performance of the proposed model (low-rank bilinear autoregression) and relevant competitors (full-rank bilinear autoregression, random forests) under different metrics: rmse (root mean squared error), mae (mean absolute error) and corr (correlation between predictions and actual observed values). Results are averaged across 24 replicated one-step-ahead forecasting exercises. The proposed low-rank bilinear autoregressive model employs $R_{A}=11$, $R_{B}=5$.
Consistent with the above goals, we study the predictive performance of the proposed model (low-rank bilinear autoregression with $R_{A}=11$, $R_{B}=5$) and two relevant competing methods. The first one (full-rank bilinear autoregression) still relies on the bilinear representation as in (1), but does not include the low-rank structure in (2). The second (random forests (Breiman, 2001)) provides instead a relevant example of a routinely implemented black-box strategy specifically designed for predictive purposes. In evaluating these competing strategies, we divide the data into a training set and a test set. The training set covers data for 204 consecutive months from $t=1$ until $t=t_{\mbox{\tiny train}}$. Leveraging these training data, we fit the three methods under analysis, including the proposed model, and then produce one-step-ahead forecasts for the log-transformed criminal activity counts at $t_{\mbox{\tiny train}}+1$ for all the 616 combinations of districts and crime categories. Recalling Section 3, under the proposed model and its full-rank version the predictions for the original log-transformed counts can be readily obtained by adding the deterministic component $\hat{\mu}_{kq(t_{\mbox{\tiny train}}+1)}$ to the forecasts for $y_{kq(t_{\mbox{\tiny train}}+1)}$ produced by these two strategies. The performance of random forests is instead assessed by training such a method directly on the observed log-counts $\log(1+c_{kqt})$, while leveraging as predictors the time index $t$, its square, the seasonal dummies ${\bf d}_{t}$ (see Section 3) and the lagged log-counts. This yields a highly flexible formulation, which can learn even more complex systematic components than those we rely for de-trending and seasonal adjustments.
Figure 3: Distribution of residual diagnostics across all crime-district series: (left) $z$-scores for mean residuals, (middle) $F$-statistics testing for 12-month seasonality, and (right) $z$-scores for linear trends. Dashed vertical lines indicate the 5% critical values ($\pm 1.96$ for mean and trend, and the 5% $F$-threshold for seasonality). The distributions are tightly centered within the thresholds, indicating satisfactory residuals behavior.
@@
The forecasts produced by the above methods are compared to the observed values of the log-transformed counts at $t_{\mbox{\tiny train}}+1$ using three different accuracy measures, namely the root mean squared error (rmse), the mean absolute deviation (mae) and the correlation (corr) between the forecasted and observed values. To provide a comprehensive assessment, these measures are further averaged over 24 replications of this predictive analysis, each obtained by shifting the training window of 204 observations forward by one month, and forecasting in the subsequent month. As a result, predictive performance is quantified on a total of $616\cdot 24=14{,}784$ point forecasts. Table 1 summarizes the outcome of this study. Interestingly, although one would expect random forests to substantially improve the forecasting accuracy of the low-rank and full-rank bilinear autoregressive models, this is not the case in such a context. In fact, both model-based solutions achieve a predictive accuracy comparable to random forests, with the low-rank formulation we propose in (1)–(2) providing essentially the same performance as its full-rank counterpart.
The above results clarify that interpretable model-based inference can be accomplished at negligible costs in predictive accuracy, and provides reassurance that the findings presented within Sections 5.2.2 and 5.2.3 are reliable and not based on an unrealistic characterization of the data-generating mechanism. This point is further underscored by Figure 3, which shows that the residuals resulting from the proposed model do not display, for all crime-district series, evidence of non-zero mean, trends or seasonal effects.
Figure 4: Estimated intra-temporal partial correlation networks between crime categories (left) and among districts (right). The colors in each network indicate clusters of nodes with high within-group absolute partial correlations and relatively low across-group dependencies (clusters are estimated via community detection algorithms (Blondel et al., 2008)). Node positions are obtained via force-direct placement methods (Fruchterman and Reingold, 1991) positioning closer in the space nodes with higher absolute partial correlations, whereas node size is proportional to the sum of absolute partial correlations involving that node. Red and blue lines denote negative and positive partial correlations, respectively, while the shading is proportional to the corresponding absolute value (higher transparency refers to values closer to 0). Finally, solid (dashed) lines refer to partial correlations whose induced posterior has $95\%$ credible intervals not including (including) the value 0.
@@
Besides achieving satisfactory predictive accuracy, a distinctive feature of the model developed in Section 34 lies in its ability to disentangle complex dependencies in the criminal activity tensor, while allowing interpretable inference on such quantities. To showcase such potentials, let us first focus on the intra-temporal dependence structures among crime categories and between districts encoded in the posteriors of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$. Figure 4 provides a graphical representation of these estimated dependencies with focus on the posterior mean of the partial correlation networks defined as $\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}\bm{\Sigma}_{A}^{-1}\mbox{diag}(\bm{\Sigma}_{A}^{-1})^{-1/2}$ and $\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}\bm{\Sigma}_{B}^{-1}\mbox{diag}(\bm{\Sigma}_{B}^{-1})^{-1/2}$ (e.g., Lauritzen, 1996). Such posterior means can be obtained by evaluating these formulas at the posterior samples of $\bm{\Sigma}_{A}$ and $\bm{\Sigma}_{B}$ produced by Algorithm 1 and then averaging across samples. Crucially, these measures provide an informative reconstruction of the graphical dependencies between crime categories and across the different districts, which quantify the intra-temporal pairwise correlations among any two generic nodes in the network conditionally on the effect of all the others, and net of the inter-temporal dependencies captured by the conditional mean in (1)–(2).
@@
Figure 4 represents graphically the aforementioned networks with nodes corresponding to crime categories and districts, respectively, and weighted edges denoting the posterior mean of the partial correlation among any pair of nodes. To facilitate interpretation and uncover relevant structures, the positions of the nodes within Figure 4 are obtained via force-direct placement methods (Fruchterman and Reingold, 1991), so that nodes with higher absolute partial correlations are displayed closer in the space. This graphical setting is also combined with the output of a community detection algorithm (Blondel et al., 2008) that unveils clusters of nodes having high within-group absolute partial correlations and relatively low across-group dependencies. As shown in the network among the different crime categories, such a perspective allows us to uncover yet unexplored modular intra-temporal dependencies between types of crimes. More specifically, the first group (green nodes) comprises, among others, crimes such as simple/aggravated battery and assault, which are physical crimes directly targeting the body. The second (purple nodes) includes, instead, crimes related to narcotics, gambling, weapons violations, and liquor license-related offenses. This clustering suggests a potential association with organized crime or coordinated gang activities in the city. The third group (blue nodes) consists mainly of property crimes such as burglary, robbery, theft and vandalism. These types of crime are often interconnected for various reasons. For example, evidence suggests that the presence of vandalism, like graffiti, can significantly increase the propensity of people to engage in property crimes, such as stealing (Keizer et al., 2008). Finally, the last group (red nodes) captures crimes which exhibit sparser dependencies, even within the community itself. This cluster encompasses crimes such as arson, homicide or fraud, which are offenses that tend to occur independently of other crimes. A partial exception in such a group is provided by prostitution, which appears in Figure 4 relatively close to the purple nodes, thereby suggesting that this category might have also direct connections with organized crime or coordinated gang activities. Besides the peculiar composition of each group, it is also worth noticing how simple battery, vandalism and narcotics play a central role within the partial correlation network, while being allocated to three different crime communities. This finding suggests that policies simultaneously targeting these three crime categories across the whole city could trigger positive spillover effects on other crime types. See Section 5.2.4 for further discussion on these results in the light of the qualitative and quantitative criminology theories presented in Section 2.
Figure 5: Left: Districts of Chicago colored according to the communities estimated from the partial correlation network in the right panel of Figure 4. Right: Plot of log-transformed distances between pairs of district centroids versus the posterior mean of the partial correlation among such district pairs (loess estimate and uncertainty bands for the relation between these two quantities are also provided).
@@
Similarly-interesting results can be obtained by examining the network structure of the partial correlations among districts displayed in the right panel of Figure 4. The community detection algorithm (Blondel et al., 2008) applied to this network reveals three distinct groups, which align with the north-east, north-west and south division of Chicago. This geographical separation becomes even more evident when visualizing the estimated group structure on a map of the city, as demonstrated in the left panel of Figure 5. The connection of the partial correlations to geographical distance is further underlined in the right panel of Figure 5, which highlights how the overall negative relationship between the pairwise distances among districts and the corresponding partial correlations is combined with heterogeneous patterns that partially depart from pure proximity structures. As discussed in Section 5.2.4, these findings contribute to the ongoing studies on criminal activities in urban environments and provide useful resources to quantify potential spillover effects across districts. Remarkably, the proposed formulation identifies spatial dependence structures without including any explicit geographical information within the model. In fact, as anticipated in Section 4.1, avoiding these constraints allows us to learn in the right panel of Figure 4 stronger dependencies within southern and north-western districts relative to those inferred among the north-eastern ones. This result partially departs from pure proximity structures that would have been imposed by, e.g., car constructions or related spatio-temporal formulations (e.g., Banerjee et al., 2015; Botella-Rocamora et al., 2015; Vicente et al., 2023), and provides quantitative evidence that supports the reported presence of organized crime and structured gangs activities in the southern and north-western area of Chicago (e.g., Aspholm, 2019; Schnell et al., 2017). This finding is confirmed by the size of the districts in Figure 4, which supports the fact that southern and north-western areas play a central role in the diffusion of crime across the city.
@@
Figure 6 displays the heatmaps of the posterior means for the elements in the matrices $\bf A$} and $\bf B$}. Recalling Section 3.1, these quantities provide insights into the “average” inter-temporal dependence among the $K$ crime categories within districts, and between the $Q$ districts for each crime category, respectively. As such, although these estimates may not directly correspond to causally-interpretable effects, Figure 6 can still unveil relevant association patterns, which complement those presented within Section 5.2.2. Recall that intra- and inter-temporal dependencies have a substantially different meaning. In fact, crime categories (or districts) displaying instantaneous intra-temporal effects, may not necessarily influence each other across different time lags.
According to Figure 6, the dominant feature of both matrices is the main diagonal, indicating a moderately high inter-temporal persistence within reports by crime categories and districts, respectively. In the posterior mean of the matrix $\bf A$} (see left panel of Figure 6) certain crimes, such as arson and embezzlement occur less frequently and in relative isolation, leading to exceptions in this overall pattern. Additionally, the occurrence of specific crime categories, such as property- and aggression-related crimes, appears to have a more remarkable significant association with the potential outbreak of specific types of crimes in the subsequent month. Such a finding provides important guidelines in the design of intervention policies guided by potential spillover effects.
Figure 6: Heatmaps of the posterior mean for the matrices $\bf A$} (left) and $\bf B$} (right) characterizing the inter-temporal dependencies under model (1)–(2). The black dots highlight parameters whose posterior distribution has $95\%$ credible intervals not including the value 0.
Relevant findings are provided also by the posterior mean of the matrix $\bf B$} in the right panel of Figure 6. Besides the direct inter-temporal persistence highlighted in the main diagonal, the estimated $\bf B$} clarifies that such a persistence acts also at a group level, especially for clusters of southern districts whose crime activities appear to be associated with those at subsequent times in other districts from the same group. Within this group it is also possible to identify districts (e.g., Chicago Lawn) with significant inter-temporal effects on the entire city. Similar impact is found also for Grand Central, situated in the north-western area. These results are in line with the spatial intra-temporal dependencies discussed in Section 5.2.2 and further support the idea that these patterns might arise from the presence of organized crime and coordinated gangs activities in the southern and north-western area of the city (Aspholm, 2019; Schnell et al., 2017).
@@
The results in Sections 5.2.25.2.3 provide novel empirical evidence in the context of criminological analyses on the mechanisms and determinants driving concentration of specific crime types in time and space (e.g., Brantingham and Brantingham, 2016, 2017; Cohen and Felson, 1979). Social, economic, and even structural characteristics of urban environments influence criminal behavior (Brantingham and Brantingham, 1995), leading to specific inter- and intra-temporal dependencies within both crime categories and spatial districts that also exhibit peculiar modular patterns (see Figure 4 and Figure 6). As such, our results underscore the importance of maintaining a systemic perspective on crime phenomena, accounting for all the three dimensions of the observed criminal activity tensor, while inferring interpretable dependencies within and across such dimensions. For example, the spatial effects we unveil in Figure 4 and Figure 6 highlight an heterogeneous distribution of crime generators and crime attractors within the urban environment (Brantingham and Brantingham, 1995; Weisburd, 2015), which, in turn, do not facilitate the same crime categories across the city. An interesting case is provided by the Chicago Loop located in the 1st district. This area serves as a financial, cultural, and commercial hub of the city, thereby increasing the number of potential targets, especially in relation to property crime. Consistent with this, the right panels of Figure 4 and Figure 6 show that such a district provides a crime generator area that showcases non-negligible intra- and inter-temporal associations with some southern and north-western districts. Recalling, Aspholm (2019) and Schnell et al. (2017), these southern and north-western districts tend to act, instead, as crime attractors due to reduced interventions and structural organizational issues that create opportunities for organized criminal activities, and hinder residents from exerting informal social control (Sampson and Groves, 1989; Shaw and McKay, 1942). These criminogenic disparities across the different areas of Chicago seem to emerge in our findings, which highlight modular intra- and inter-temporal associations among Chicago districts, with southern and north-western areas displaying more dense patterns.
@@
The results in Figure 5 further underscore the importance of the spatial proximity and potential spillover effects. According to our findings, there is an overall negative relationship between the distance among districts and the corresponding partial correlation. A possible explanation involves mechanisms that combine residential location and social interactions (Zenou, 2003), highlighting the role of behavior in disadvantaged environments (Glaeser et al., 1996). For instance, evidence suggests that offenders are inclined to exploit criminal opportunities close to their residence (Bernasco, 2010; Bernasco and Block, 2009; Bernasco et al., 2014; Hipp, 2016; Menting et al., 2016; Townsley et al., 2016). Consequently, the concentration of offenders in a neighborhood can trigger contagion effects, reinforcing negative social dynamics (Becker and Murphy, 2009).
@@
Moreover, increased criminal concentration intensifies the competition among offenders, reducing the expected gain of criminal activity (Becker, 1968) and encouraging repeat offenses (Zenou, 2003). This is also consistent with research on repeat victimization and temporal clustering, which indicates that the occurrence of an initial crime increases the probability of subsequent offenses in the same location. Offenders tend to exploit previously acquired knowledge about susceptible targets and the absence of effective interventions (e.g., Johnson and Bowers, 2004; Townsley et al., 2003). This persistence of crime is also evident in our findings (see Figure 6), although there are differences across categories. Specifically, while the main diagonal of the left matrix in Figure 6 shows a clear temporal persistence for property- and aggression-related crimes, as well as for organized crime offenses like narcotics, prostitution, and gambling, this pattern is less pronounced for violent offenses such as sexual-related crimes, homicides and manslaughter, and for property crimes like arson and embezzlement. These inferred heterogeneities in crime persistence by offense type provide novel empirical evidence to the literature on such phenomena (Chainey, 2021; Haberman and Ratcliffe, 2012; Johnson and Bowers, 2004; Johnson et al., 2009).